Hi! I need help with this direct variation equation.
It says If y=4 when x=2, find y when x=16.
How do you solve this?
Thanks -Drea

Found 2 solutions by nerdybill, richwmiller:
Answer by nerdybill(7384)

It says If y=4 when x=2, find y when x=16.
How do you solve this?
.
In "direct variation" the relation is:
y = kx
where
k is a constant
.
The problem gives you the fact that:
If y=4 when x=2
Use this to discover what k is:
Plugging what we know into:
y = kx
4 = k(2)
4/2 = k
2 = k
.
So, our relation is:
y = 2x
To find y when x=16 simply plug x into:
y = 2x
y = 2(16)
y = 32

Answer by richwmiller(17219)


I looked this up and they don't consider it direct variation if b does not equal zero so b always has to equal zero for direct variation
y=mx is the equation that governs direct variation.
4=m*2
If m=2
y=2*16
y=32